OpenStudy (megannicole51):

Find the Taylor polynomial of degree n=4 for x near the point a=pi/2 for the function cos(2x) P4(x)=

OpenStudy (megannicole51):

i dont understand what to do with the a....

OpenStudy (megannicole51):

@agent0smith

OpenStudy (agent0smith):

http://tutorial.math.lamar.edu/Classes/CalcII/TaylorSeries_files/eq0021M.gif a is that a. the other stuff is just the value of the function at a, and it's various derivatives It's much more complex than it looks, though.

OpenStudy (megannicole51):

so what do i do? :/

OpenStudy (agent0smith):

Find f(a) which is f(pi/2) find the first derivative, then find f'(pi/2) find the second... etc down to the fourth derivative

OpenStudy (agent0smith):

eg f(a) = cos(2*pi/2) f' = -2sin(2x) f'(a) = -2sin(2pi/2)

OpenStudy (agent0smith):

Oh you only need the third derivative, for n=4.

OpenStudy (anonymous):

cos(2a) = cos(2*pi/2) = cos(pi) = -1. sin(2a) = sin(pi) = 0. The taylor series of cos(2x) at a=2pi is: \[\cos(\pi)-\frac{ 2\sin(\pi) }{ 1!}(x-2\pi)+\frac{ 4\cos(\pi) }{ 2! }(x-2\pi)^2-\frac{ 8\sin(\pi) }{ 3! }(x-2\pi)^3+\frac{ 16\cos(\pi) }{ 4! }(x-2\pi)^4\]

OpenStudy (megannicole51):

a=pi/2

OpenStudy (anonymous):

\[= -1 -2(x-2\pi)^2 -\frac{ 2 }{ 3 }(x-2\pi)^4\]

OpenStudy (megannicole51):

it says its wrong:/

OpenStudy (anonymous):

Woops, It's pi/2 instead of 2pi and the symbols are wrong. Sorry. D:

OpenStudy (megannicole51):

no its okay!

OpenStudy (anonymous):

Give me a second, though.

OpenStudy (anonymous):

Okay. The Taylor polynomial of degree four near point a is: \[p(x) = f(a) + \frac{ f'(a)}{ 1! }(x-a)+\frac{ f''(a) }{ 2! }(x-a)^2+\frac{ f'''(a) }{ 3! }(x-a)^3 +\frac{ f''''(a) }{ 4! }(x-a)^4\] So, be f(x)=cos(2x) and a =pi/2 \[f(a) = \cos(\frac{ 2\pi }{ 2 }) = \cos(\pi) = -1\] \[f'(a) = -2\sin(2a) = -2\sin(\frac{ 2\pi }{ 2 }) = -2 \sin(\pi) = 0\] \[f''(a)=-4\cos(2a) = -4\cos(\pi) = 4\] \[f'''(a) = 8\sin(2a)=8\sin(\pi)=0\] \[f''''(a)=16\cos(2a) = 16\cos(\pi)=-16\]

OpenStudy (megannicole51):

its not right

OpenStudy (anonymous):

Substituting this values in the taylor polynomial we get: \[p(x) = -1 + \frac{ 0 }{ 1! }(x-a)+\frac{ 4 }{ 2! }(x-a)^2+\frac{ 0 }{ 3! }(x-a)^3-\frac{ 16 }{ 4! }(x-a)^4\] \[= -1 +2(x-a)^2-\frac{ 2 }{ 3 }(x-a)^4\] Remembering that a = pi/2 .... \[P(x) = -1+2(x-\frac{ \pi }{ 2 })^2-\frac{ 2 }{ 3 }(x-\frac{ \pi }{ 2 })^4\]

OpenStudy (megannicole51):

ooooh okay yeah its right! thank you:)

OpenStudy (anonymous):

You just have to calculate f(a), f'(a), f''(a), f"'(a) and f""(a), plug those in the taylor polynomial P(x) and substitute where necessary. It's a real chore. D: Just remember to get the positive/negative sign correctly when dividing/subs.